OpenStudy (anonymous):

Who ever helps me ill give a medal So please help [6.02] For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

If you isolate x, you have -x=2y-4 --->x=-2y+4

OpenStudy (anonymous):

thank u can help me with some more

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely Many Solutions No Solutions

OpenStudy (anonymous):

multiply the first equation by two and you get: 6x-4y=12 Now you have two equations one says 6x-4y=12 and the other one says: 6x-4y=14 As left part of both equations is the same (6x-4y), then right parts should be the same, however 12 is not equal to 14. That means that the system has no solutions

OpenStudy (anonymous):

u make easy to understand

OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -x + 2 (1, 6) (-2, -1) (-3, 2) (1, 1)

OpenStudy (anonymous):

Any point that fulfills y-x>4 and y+x>2 will also fulfills (adding both inequalities) 2y>6--->y>3 the pair (1,6) is the solution